OpenStudy (anonymous):

A car traveling initially at 8.9 m/s accelerates at the rate of 0.875 m/s^2 for 2.9s What is its velocity at the end of the acceleration?

OpenStudy (anonymous):

V= V(initial) + at V(initial)= 8.9 m/s a=0.875 m/s^2 t=2.9s V= ? V= 8.9 + (0.875)(2.9) V= 8.9+ 2.5375 V= 11.4375m/s

OpenStudy (anonymous):

\[V=V_0+a·t=8.9+0.875·2.9=11.44 m/s\]

OpenStudy (anonymous):

Thanks for the help :)

OpenStudy (anonymous):

you are welcome